Q2: Find the general solution to T = AT where A is given by 1-3 -2] A = 2 1] and then sketch a variety of solution curves.
Answer #1

Here on ZAR where A = [-3 -2] we can rewrite the system as and de 1 2 11=-3+2 de 2-3x, -20 e for equilibrium point is given bNext step is to find 7 from ① and w as 22-1 we need to solve 5) -241-28291 » 20,+ 2.822-1 » Batdzak = B22-d2t when drak, B2 =

the solution curve is singular node or stable singular node.